# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 2.4 Transition Dynamics
In conventional PBRFT, the transition dynamics are deterministic: the next state is determined once an action is made, as follows:
@@
Text actions $(\mathcal{A}_{\text{text}})$ generate natural language outputs without altering the environment state. Structured actions $(\mathcal{A}_{\text{action}})$, delimited by <action_start> and <action_end>, can either query external tools or directly modify the environment. This sequential formulation contrasts with the one-shot mapping of PBRFT, enabling policies that iteratively combine communication, information acquisition, and environment manipulation.
@@ 2.5 Reward Function
PBRFT commonly features a reward function with verifiable response correctness, which may be implemented using either a rule-based verifier [31] or a neural network-parameterized reward model [34]. Regardless of the implementation approach, its core follows the equation:
$\mathcal{R}_{\text{trad}}(s_{0},a)=r(a),$ (8)
where $r:\mathcal{A}\!\to\!\mathbb{R}$ is a scalar score supplied by a human- or AI-preference model; with no intermediate feedback.
@@
$\mathcal{R}_{\text{agent}}(s_{t},a_{t})=\begin{cases}r_{\text{task}}&\text{on task completion},\\[2.0pt] r_{\text{sub}}(s_{t},a_{t})&\text{for step-level progress},\\[2.0pt] 0&\text{otherwise},\end{cases}$ (9)
allowing dense, sparse, or learned rewards (e.g., unit-test pass, symbolic verifier success).
@@ 2.6 Learning Objective
$J_{\text{agent}}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\,\sum_{t=0}^{T-1}\gamma^{t}R_{\text{agent}}(s_{t},a_{t})\right],\qquad 0<\gamma<1,$ (11)
optimized via policy-gradient or value-based methods with exploration and long-term credit assignment.
PBRFT focuses on single-turn text quality alignment without explicit planning, tool use, or environment feedback, while agentic RL involves multi-turn planning, adaptive tool invocation, stateful memory, and long-horizon credit assignment, enabling the LLM to function as an autonomous decision-making agent.
@@ 2.7 RL Algorithms
In contemporary research, RL algorithms constitute a pivotal component in both PBRFT and agentic RL frameworks. Different RL algorithms demonstrate distinct sample efficiency and performance characteristics, each offering a unique approach to the central challenge of aligning model outputs with complex, often subjective, human goals. The canonical methods, such as REINFORCE, PPO [12], GRPO [31], and DPO [29], form a spectrum from general policy gradients to specialized preference learning. We next introduce each of these three classic algorithms and provide a comparison of popular variants from each family in Table 2.
@@
As one of the earliest policy gradient algorithms, REINFORCE provides the foundational theory for training stochastic policies. It operates by increasing the probability of actions that lead to high cumulative reward and decreasing the probability of those that lead to low reward. Its objective function is given by:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{s_{0}}\left[\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{R}(s_{0},a^{(i)})-b(s_{0})\right)\nabla_{\theta}\log\pi_{\theta}(a^{(i)}|s_{0})\right],$ (12)
where $a^{(i)}\sim\pi_{\theta}(a|s_{0})$ is the $i$-th sampled response, $\mathcal{R}(s_{0},a)$ denotes the final rewards received on task completion, and $b(s)$ is a baseline function to reduce the variance of the policy gradient estimate. In general, $b(s)$ can be any function, including random variables.
@@
PPO [12] became the dominant RL algorithm for LLM alignment due to its stability and reliability. It improves upon vanilla policy gradients by limiting the update step to prevent destructively large policy changes. Its primary clipped objective function is:
$L_{PPO}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\min\left(\frac{\pi_{\theta}(a_{t}^{(i)}|s_{t})}{\pi_{\theta_{old}}(a_{t}^{(i)}|s_{t})}A(s_{t},a_{t}^{(i)}),\;\;\mathrm{clip}\left(\frac{\pi_{\theta}(a_{t}^{(i)}|s_{t})}{\pi_{\theta_{old}}(a_{t}^{(i)}|s_{t})},1-\epsilon,1+\epsilon\right)A(s_{t},a_{t}^{(i)})\right),$ (13)
@@
$A(s_{t},a_{t})=\mathcal{R}(s_{t},a_{t})-V(s_{t}),$ (14)
where $V_{\theta}(s)$ is the learned value function, i.e., the expectation $\mathbb{E}_{a\sim\pi_{\theta}(a|s)}[\mathcal{R}(s,a)]$, which is derived from a critic network that is of the same size as the policy network. The clip term prevents the probability ratio from moving too far from 1, ensuring stable updates. A key drawback is its reliance on a separate critic network for advantage estimation, which substantially increases the parameter count during training.
@@
DPO represents a groundbreaking shift by entirely bypassing the need for a separate reward model. It reframes the problem of maximizing a reward under a KL-constraint as a likelihood-based objective on human preference data. Given a dataset of preferences $D=\{(y_{w},y_{l})\}$, where $y_{w}$ is the preferred response and $y_{l}$ is the dispreferred one, the DPO loss is:
$L_{DPO}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}\right)\right],$ (15)
where $\pi_{ref}$ is a reference policy (usually the initial SFT model), and $\beta$ is a hyperparameter. While DPO eliminates the critic, its performance is intrinsically tied to the quality and coverage of its static preference dataset. Variants have emerged to address its limitations, including IPO (Identity Preference Optimization) [35] which adds a regularization term to prevent overfitting, and KTO (Kahneman-Tversky Optimization) [36], which learns from per-response binary signals (desirable/undesirable) rather than strict pairwise comparisons. See Table. 2 for more variants.
@@
The remarkable success achieved by DeepSeek has catalyzed significant research interest in GRPO. Proposed to address the inefficiency of PPO’s large critic, GRPO introduces a novel, lightweight evaluation paradigm. It operates on groups of responses, using their relative rewards within a group to compute advantages, thus eliminating the need for an absolute value critic. The core GRPO objective can be conceptualized as:
$L_{GRPO}=\frac{1}{G}\sum_{g=1}^{G}\min\left(\frac{\pi_{\theta}(a_{t}^{(g)}|s_{t}^{(g)})}{\pi_{\theta_{old}}(a_{t}^{(g)}|s_{t}^{(g)})}\hat{A}(s_{t}^{(g)},a_{t}^{(g)}),\;\;\mathrm{clip}\left(\frac{\pi_{\theta}(a_{t}^{(g)}|s_{t}^{(g)})}{\pi_{\theta_{old}}(a_{t}^{(g)}|s_{t}^{(g)})},1-\epsilon,1+\epsilon\right)\hat{A}(s_{t}^{(g)},a_{t}^{(g)})\right),$ (16)
@@
This group-relative approach is highly sample-efficient and reduces computational overhead. Consequently, a series of novel algorithms derived from the GRPO framework have been subsequently proposed (see Table. 2), aiming to substantially enhance both the sample efficiency and asymptotic performance of reinforcement learning methodologies.
@@
Method Year Objective Type Clip KL Penalty Key Mechanism Signal
@@
PPO [12] 2017 Policy gradient Yes No Policy ratio clipping Reward
VAPO [37] 2025 Policy gradient Yes Adaptive Adaptive KL penalty + variance control Reward + variance signal
PF-PPO [38] 2024 Policy gradient Yes Yes Policy filtration Noisy reward
VinePPO [39] 2024 Policy gradient Yes Yes Unbiased value estimates Reward
PSGPO [40] 2024 Policy gradient Yes Yes Process supervision Process Reward
@@
DPO [29] 2024 Preference optimization No Yes Implicit reward related to the policy Human preference
$\beta$-DPO [41] 2024 Preference optimization No Adaptive Dynamic KL coefficient Human preference
SimPO [42] 2024 Preference optimization No Scaled Use the average log probability of a sequence as the implicit reward Human preference
IPO [35] 2024 Implicit preference No No Leverage generative LLMs as preference classifiers for reducing the dependence on external human feedback or reward models Preference rank
KTO [36] 2024 Knowledge transfer optimization No Yes Teacher stabilization Teacher-student logit
ORPO [43] 2024 Online regularized
@@
No Yes Online stabilization Online feedback reward
Step-DPO [44] 2024 Preference optimization No Yes Step-wise supervision Step-wise preference
LCPO [45] 2025 Preference optimization No Yes Length preference with
limited data and training
Reward
@@
GRPO [31] 2025 Policy Gradient
@@
Yes Yes Group-based relative reward
@@
Group-based reward
DAPO [46] 2025 Surrogate of GRPO’s Yes Yes Decoupled clip and dynamic sampling Dynamic group-based reward
GSPO [47] 2025 Surrogate of GRPO’s Yes Yes Define the importance ratio based on sequence likelihood and performs sequence-level clipping, rewarding, and optimization Smooth group-based reward
GMPO [48] 2025 Surrogate of GRPO’s Yes Yes Geometric mean of token-level rewards Margin-based reward
ProRL [49] 2025 Same as GRPO’s Yes Yes Reference policy reset Group-based reward
Posterior-GRPO [50] 2025 Same as GRPO’s Yes Yes Reward only successful processes Process-based reward
Dr.GRPO [51] 2025 Unbiased GRPO’s objective Yes Yes Eliminate the bias in
@@
Group-based reward
Step-GRPO [52] 2025 Same as GRPO’s Yes Yes Rule-based reasoning rewards Step-wise reward
SRPO [53] 2025 Same as GRPO’s Yes Yes Two-staged history-resampling Reward
GRESO [54] 2025 Same as GRPO’s Yes Yes Pre-rollout filtering Reward
StarPO [55] 2025 Same as GRPO’s Yes Yes Reasoning-guided actions for
@@
Group-based reward
GHPO [56] 2025 Policy gradient Yes Yes Adaptive prompt refinement Reward
Skywork R1V2 [57] 2025 GRPO’s with hybrid reward signal Yes Yes Selective sample buffer Multimodal reward
ASPO [58] 2025 GRPO’s with shaped advantage function Yes Yes Apply a clipped bias directly to advantage function Group-based reward
TreePo [59] 2025 Same as GRPO’s Yes Yes Self-guided policy rollout for reducing the compute burden Group-based reward
EDGE-GRPO [60] 2025 Same as GRPO’s Yes Yes Entropy-driven advantage and duided error correction to mitigate advantage collapse Group-based reward
DARS [61] 2025 Same as GRPO’s Yes No Reallocate compute from medium-difficulty to the hardest problems via multi-stage rollout sampling Group-based reward
CHORD [62] 2025 Weighted sum of GRPO’s and Supervised Fine-Tuning losses Yes Yes Reframe Supervised Fine-Tuning as a dynamically weighted auxiliary objective within the on-policy RL process Group-based reward
PAPO [63] 2025 Surrogate of GRPO’s Yes Yes Encourage learning to perceive while learning to reason through the Implicit Perception Loss Group-based reward
Pass@k Training 2025 Same as GRPO’s Yes Yes Pass@k metric as the reward to continually train a model Group-based reward
@@ 3 Agentic RL: The model capability perspective
In this section, we conceptually characterize Agentic RL as the principled training of an autonomous agent composed of a set of key abilities/modules, i.e., planning (Section 3.1), tool use (Section 3.2), memory (Section 3.3), self-improvement (Section 3.4), reasoning (Section 3.5), perception (Section 3.6), and others (Section 3.7), following the classic LLM agent definition [64, 65], as demonstrated in Figure˜4. Traditionally, an agent pairs an LLM with mechanisms for planning (e.g., task decomposition and plan selection) [66], reasoning (chain-of-thought or multi-turn inference) [67], external tool invocation [68], long- and short-term memory, and iterative reflection to self-correct and refine behavior. Agentic RL thus treats these components not as static pipelines but as interdependent policies that can be jointly optimized: RL for planning learns multi-step decision trajectories; RL for memory shapes retrieval and encoding dynamics; RL for tool use optimizes invocation timing and fidelity; and RL for reflection drives internal self-supervision and self-improvement. Consequently, our survey systematically examines how RL empowers planning, tool use, memory, reflection, and reasoning in subsequent subsections. We aim to provide a high-level conceptual delineation of RL’s applications for agent capabilities, rather than an exhaustive enumeration of all related work, which we provide in Section 6.
Figure 3: The dynamic interaction process between agentic LLMs and the environment.
Figure 4: A summary of the overall six aspects where RL empowers agentic LLMs. Note that the representative methods listed here are not exhaustive; please refer to our main text.

### v4

@@ 2.4 Transition Dynamics
In conventional PBRFT, the transition dynamics are deterministic: the next state is determined once an action is taken, as follows:
@@
Text actions $(\mathcal{A}_{\text{text}})$ generate natural language outputs without altering the environmental state. Structured actions $(\mathcal{A}_{\text{action}})$, delimited by <action_start> and <action_end>, can either query external tools or directly modify the environment. This sequential formulation contrasts with the one-shot mapping of PBRFT, enabling policies that iteratively combine communication, information acquisition, and environment manipulation.
@@ 2.5 Reward Function
PBRFT commonly features a reward function with verifiable response correctness, which may be implemented using either a rule-based verifier (DeepSeek-AI et al., 2025) or a neural network-parameterized reward model (Zhong et al., 2025). Regardless of the implementation approach, its core follows the equation:
$\mathcal{R}_{\text{trad}}(s_{0},a)=r(a).$ (8)
where $r:\mathcal{A}\!\to\!\mathbb{R}$ is a scalar score supplied by a human- or AI-preference model, with no intermediate feedback.
@@
$\mathcal{R}_{\text{agent}}(s_{t},a_{t})=\begin{cases}r_{\text{task}}&\text{on task completion},\\[2.0pt] r_{\text{sub}}(s_{t},a_{t})&\text{for step-level progress},\\[2.0pt] 0&\text{otherwise}.\end{cases}$ (9)
allowing dense, sparse, or learned rewards (e.g., unit-test passes, symbolic verifier success).
@@ 2.6 Learning Objective
$J_{\text{agent}}(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\,\sum_{t=0}^{T-1}\gamma^{t}R_{\text{agent}}(s_{t},a_{t})\right],\qquad 0<\gamma<1.$ (11)
This objective is optimized via policy‑gradient or value‑based methods with exploration and long‑term credit assignment.
PBRFT focuses on single-turn text quality alignment without explicit planning, tool use, or environmental feedback, while Agentic RL involves multi-turn planning, adaptive tool invocation, stateful memory, and long‑horizon credit assignment, enabling the LLM to function as an autonomous decision‑making agent.
@@ 2.7 RL Algorithms
In contemporary research, RL algorithms constitute a pivotal component in both PBRFT and Agentic RL frameworks. Different RL algorithms demonstrate distinct sample efficiency and performance characteristics, each offering a unique approach to the central challenge of aligning model outputs with complex, often subjective, human goals. The canonical methods, such as REINFORCE, PPO (Schulman et al., 2017), GRPO (DeepSeek-AI et al., 2025), and DPO (Rafailov et al., 2023), form a spectrum from general policy gradients to specialized preference learning. We next introduce each of these four classic algorithms and provide a comparison of popular variants from each family in Table 2.
@@
As one of the earliest policy gradient algorithms, REINFORCE (Williams, 1992) provides the foundational theory for training stochastic policies. It operates by increasing the probability of actions that lead to high cumulative reward and decreasing the probability of those that lead to low reward. Its objective function is given by:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{s_{0}}\left[\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{R}(s_{0},a^{(i)})-b(s_{0})\right)\nabla_{\theta}\log\pi_{\theta}(a^{(i)}|s_{0})\right].$ (12)
where $a^{(i)}\sim\pi_{\theta}(a|s_{0})$ is the $i$-th sampled response, $\mathcal{R}(s_{0},a)$ denotes the final rewards received on task completion, and $b(s)$ is a baseline function to reduce the variance of the policy gradient estimate. In general, $b(s)$ can be any function, including random variables. In practice, $b(s)$ is commonly instantiated as the value function $V(s)$. Despite with advantages of the concise formula and easy implementation, REINFORCE suffers from drawbacks such as high variance in gradient estimates, sample inefficiency, sensitivity to learning rate and the lack of a critic (value estimator).
@@
PPO (Schulman et al., 2017) became the dominant RL algorithm for LLM alignment due to its stability and reliability. It improves upon vanilla policy gradients by limiting the update step to prevent destructively large policy changes. Its primary clipped objective function is:
$L_{PPO}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\min\left(\frac{\pi_{\theta}(a_{t}^{(i)}|s_{t})}{\pi_{\theta_{old}}(a_{t}^{(i)}|s_{t})}A(s_{t},a_{t}^{(i)}),\;\;\mathrm{clip}\left(\frac{\pi_{\theta}(a_{t}^{(i)}|s_{t})}{\pi_{\theta_{old}}(a_{t}^{(i)}|s_{t})},1-\epsilon,1+\epsilon\right)A(s_{t},a_{t}^{(i)})\right).$ (13)
@@
$A(s_{t},a_{t})=\mathcal{R}(s_{t},a_{t})-V(s_{t}).$ (14)
where $V_{\theta}(s)$ is the learned value function, i.e., the expectation $\mathbb{E}_{a\sim\pi_{\theta}(a|s)}[\mathcal{R}(s,a)]$, which is typically, but not necessarily, derived from a critic network that is of the same size as the policy network. The clip term prevents the probability ratio from moving too far from 1, ensuring stable updates. The estimation of the advantage function plays a predominant role in the performance of PPO. Recent variants have concentrated on reducing the bias (Kazemnejad et al., 2024) or variance (Yue et al., 2025b) in the advantage estimation. Meanwhile, some other variants make improvements from the perspectives of stable policy update mechanisms (Liu et al., 2025s) or mitigating sparse rewards (Dai et al., 2025). Despite these improvements, a remaining drawback is its reliance on a separate critic network for advantage estimation, which substantially increases the parameter count during training.
@@
DPO represents a groundbreaking shift by entirely bypassing the need for a separate explicit reward model. It reframes the problem of maximizing a reward under a KL-constraint as a likelihood-based objective on human preference data. Given a dataset of preferences $D=\{(y_{w},y_{l})\}$, where $y_{w}$ is the preferred response and $y_{l}$ is the dispreferred one, the DPO loss is:
$L_{DPO}(\pi_{\theta};\pi_{ref})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)}\right)\right].$ (15)
where $\pi_{ref}$ is a reference policy (usually the initial SFT model), and $\beta$ is a hyperparameter. While DPO eliminates the critic, its performance is intrinsically tied to the quality and coverage of its static preference dataset. Variants have emerged to address its limitations via involving external or online data (Ethayarajh et al., 2024; Hong et al., 2024a). In addition, some other work attempts to improve by introducing generalized optimization objectives (Gheshlaghi Azar et al., 2024) or sophisticated implicit reward mechanisms (Meng et al., 2024; Lai et al., 2024; Hong et al., 2025a).
@@
The remarkable success achieved by DeepSeek (Guo et al., 2025a) has catalyzed significant research interest in GRPO. Proposed to address the inefficiency of PPO’s large critic, GRPO introduces a novel, lightweight evaluation paradigm. It operates on groups of responses, using their relative rewards within a group to compute advantages, thus eliminating the need for an absolute value critic. The core GRPO objective can be conceptualized as:
$L_{GRPO}=\frac{1}{G}\sum_{g=1}^{G}\min\left(\frac{\pi_{\theta}(a_{t}^{(g)}|s_{t}^{(g)})}{\pi_{\theta_{old}}(a_{t}^{(g)}|s_{t}^{(g)})}\hat{A}(s_{t}^{(g)},a_{t}^{(g)}),\;\;\mathrm{clip}\left(\frac{\pi_{\theta}(a_{t}^{(g)}|s_{t}^{(g)})}{\pi_{\theta_{old}}(a_{t}^{(g)}|s_{t}^{(g)})},1-\epsilon,1+\epsilon\right)\hat{A}(s_{t}^{(g)},a_{t}^{(g)})\right).$ (16)
@@
This group-relative approach is highly sample-efficient and reduces computational overhead. However, the group-based advantage estimation is vulnerable to high variance and low accuracy. Consequently, a series of novel algorithms derived from the GRPO framework have been subsequently proposed (see Table 2), aiming to substantially improve its advantage estimation.
@@
Method Objective Type Key Mechanism
@@
PPO (Schulman et al., 2017) Policy gradient Policy ratio clipping
VAPO (Yue et al., 2025b) Policy gradient Adaptive KL penalty + variance control
LitePPO (Liu et al., 2025s) Policy gradient Stable advantage updates
PF-PPO (Zhang et al., 2025c) Policy gradient Policy filtration
VinePPO (Kazemnejad et al., 2024) Policy gradient Unbiased value estimates
PSGPO (Dai et al., 2025) Policy gradient Process supervision
@@
DPO (Rafailov et al., 2023) Preference optimization Implicit reward related to the policy
$\beta$-DPO (Wu et al., 2024) Preference optimization Dynamic KL coefficient
SimPO (Meng et al., 2024) Preference optimization Use the average log probability of a sequence as the implicit reward
IPO (Gheshlaghi Azar et al., 2024) A special case of a more general objective exclusively expressed in terms of pairwise preferences Always regularizes its solution towards a preference policy by controlling the gap between the log-likelihood ratios, which avoids the over-fitting to the preference dataset.
KTO (Ethayarajh et al., 2024) Knowledge transfer optimization Teacher stabilization
ORPO (Hong et al., 2024a) Online regularized
@@
Online stabilization
Step-DPO (Lai et al., 2024) Preference optimization Step-wise supervision
LCPO (Hong et al., 2025a) Preference optimization Length preference with limited data and training
@@
GRPO (DeepSeek-AI et al., 2025) Policy Gradient
@@
Group-based relative reward
@@
DAPO (Yu et al., 2025e) Surrogate of GRPO’s Decoupled clip and dynamic sampling
GSPO (Zheng et al., 2025a) Surrogate of GRPO’s Define the importance ratio based on sequence likelihood and performs sequence-level clipping, rewarding, and optimization
GMPO (Zhao et al., 2025f) Surrogate of GRPO’s Geometric mean of token-level rewards
ProRL (Liu et al., 2025h) Same as GRPO’s Reference policy reset
Posterior-GRPO (Fan et al., 2025a) Same as GRPO’s Reward only successful processes
Dr.GRPO (Liu et al., 2025r) Unbiased GRPO’s objective Eliminate the bias in
@@
Step-GRPO (Zhang et al., 2025j) Same as GRPO’s Rule-based reasoning rewards
SRPO (Zhang et al., 2025s) Same as GRPO’s Two-staged history-resampling
GRESO (Zheng et al., 2025b) Same as GRPO’s Pre-rollout filtering
StarPO (Wang et al., 2025v) Same as GRPO’s Reasoning-guided actions for
@@
GHPO (Liu et al., 2025u) Policy gradient Adaptive prompt refinement
Skywork R1V2 (Wang et al., 2025i) GRPO’s with hybrid reward signal Selective sample buffer
ASPO (Lin & Xu, 2025) GRPO’s with shaped advantage function Apply a clipped bias directly to advantage function
TreePo (Li et al., 2025n) Same as GRPO’s Self-guided policy rollout for reducing the compute burden
EDGE-GRPO (Zhang et al., 2025c) Same as GRPO’s Entropy-driven advantage and duided error correction to mitigate advantage collapse
DARS (Yang et al., 2025h) Same as GRPO’s Reallocate compute from medium-difficulty to the hardest problems via multi-stage rollout sampling
CHORD (Zhang et al., 2025q) Weighted sum of GRPO’s and Supervised Fine-Tuning losses Reframe Supervised Fine-Tuning as a dynamically weighted auxiliary objective within the on-policy RL process
PAPO (Wang et al., 2025u) Surrogate of GRPO’s Encourage learning to perceive while learning to reason through the Implicit Perception Loss
Pass@k Training (Chen et al., 2025l) Same as GRPO’s Pass@k metric as the reward to continually train a model
@@ 3 Agentic RL: The model capability perspective
Figure 3: The agent–environment interaction and RL loop for agentic LLMs. Core agentic capabilities drive action generation, while the environment provides feedback and rewards, which are aggregated through RL-based optimization across diverse task domains (“Collab.” denotes tasks requiring explicit task division and multi-agent coordination).
In this section, we conceptually characterize Agentic RL as the principled training of an autonomous agent composed of a set of key abilities/modules, i.e., planning (Section 3.1), tool use (Section 3.2), memory (Section 3.3), self-improvement (Section 3.4), reasoning (Section 3.5), perception (Section 3.6), and others (Section 3.7), following the classic LLM agent definition (Weng, 2023; Shang et al., 2025b), as demonstrated in Figure˜5. Traditionally, an agent pairs an LLM with mechanisms for planning (e.g., task decomposition and plan selection) (Wei et al., 2025a), reasoning (chain-of-thought or multi-turn inference) (Zhang et al., 2024c), external tool invocation (Qin et al., 2024b), long- and short-term memory, and iterative reflection to self-correct and refine behavior. Agentic RL thus treats these components not as static pipelines but as interdependent policies that can be jointly optimized: RL for planning learns multi-step decision trajectories; RL for memory shapes retrieval and encoding dynamics; RL for tool use optimizes invocation timing and fidelity; and RL for reflection drives internal self‑supervision and self-improvement. Consequently, our survey systematically examines how RL empowers planning, tool use, memory, reflection, and reasoning in subsequent subsections. We aim to provide a high-level conceptual delineation of RL’s applications for agent capabilities, rather than an exhaustive enumeration of all related work, which we provide in Section 6.